OpenStudy (anonymous):

rearrange this formular to make r the subject. A=4πr²

OpenStudy (saifoo.khan):

sqrt(A/4π)

OpenStudy (anonymous):

sqrt? I don't understand

OpenStudy (saifoo.khan):

\[\sqrt{A/4π} \]

OpenStudy (anonymous):

so, r=√A/4π?

OpenStudy (saifoo.khan):

yes.

myininaya (myininaya):

\[A=4 \pi r^2\] \[r^2=A/(4 \pi)\] take square root of both sides

OpenStudy (anonymous):

how did you work that out? It's for an upcoming exam

OpenStudy (saifoo.khan):

look, 4π is being multiplied on the right... so it will b divided with A. A/4π = r^2

OpenStudy (saifoo.khan):

then a Sqrt to cancel the square.

OpenStudy (saifoo.khan):

finally u r done with r.

OpenStudy (anonymous):

thanks c:

OpenStudy (saifoo.khan):

C: